# Decentralized Droop-based Finite-Control-Set Model Predictive Control of Inverter-based Resources in Islanded AC Microgrid

### v1

@@
Abstract
@@
keywords:
Droop control, finite-control-set model predictive control, active power sharing, reactive power sharing, voltage regulation, inverter-based resources, islanded microgrid.
@@ 1 Introduction
Traditional Microgrids (MGs) are a conglomeration of numerous fossil-fuel-driven power generation elements like synchronous machines (SMs). However, in recent times, an increase in penetration of inverter-based resources (IBRs) such as solar and wind generation has caused a shift in power generation trends in the existing infrastructure. The addition of energy storage elements (ESEs) further upped the choice of power allocation in the MGs in terms of the available degree of freedom. Based on their configuration and the mode of operation MGs are classified into two categories known as grid-connected and islanded modes Hatziargyriou et al. (2007). Control of the traditional MG employs a hierarchical control structure Guerrero et al. (2011), with droop-control being the lowest level followed by the power and the energy management (EM) layers. The impending integration of IBRs into the traditional MGs is giving rise to structural and operational changes in the form of transitioning away from the SMs which are ramp-limited due to their inertial properties. However, the low inertia and the fast switching properties of the IBRs pose a major challenge in the form of assessing their stability and robustness Milano et al. (2018). One such method that is widely embraced to restore the inertia to the grid is to introduce virtual inertia Anubi and Ameli (2022); Cady et al. (2017).
@@
Regardless of the inertial properties of the SMs and the IBRs, the control objective remains unequivocal in terms of balancing the active and reactive power in the MGs while simultaneously maintaining the grid frequency around the desired operating point. Droop control is the method widely employed in addressing the control of the power and the frequency namely voltage droop and frequency droop Silva et al. (2021). The IBRs are operated in parallel mode due to high reliability and flexibility in terms of operation. However, the application of droop control in such a parallel setting is not suitable while sharing a nonlinear load due to the presence of harmonic currents Khaledian and Golkar (2017). Moreover, the frequency and amplitude deviation in the IBRs droop settings are dependent on the load scenarios Guerrero et al. (2013). Addition of virtual control loops such as adding a virtual resistance into the droop methods and decentralizing the control (consensus-based) are proposed to address the complications arising from the presence of current harmonics Lu et al. (2018). Although Communication-based and consensus control algorithms offer efficient voltage regulation and effective power sharing among connected IBRs, the methods suffer from the huge cost and reliability of communication links. On the other hand, droop control does not require communication between IBRs to effect power sharing and voltage regulation Dong et al. (2022). Multiple loops of conventional droop techniques often result in poor power-sharing due to parameter variations, especially in the current loop that generates PWM signals for the inverters He et al. (2012).
@@
Thus, in this paper, we utilize FCS-MPC in the inner loop to directly generate optimal switching signals for the IBRs. Finite-control-set model predictive control (FCS-MPC) has emerged as a promising discrete-time model-based nonlinear predictive control algorithm applicable to fast-switching power electronics dominated systems. Unlike the continuous-set MPC that requires modulators and a longer horizon to generate the pulse-width modulated (PWM) signals, FCS-MPC eliminates the need for modulators and accounts for nonlinearities in the power converter dynamics Bayhan et al. (2016). It uses the horizon of h = 1, which significantly reduces the computational burden that may result from very low sampling time in high-frequency converters. To this end, we develop a decentralized droop-based FCS-MPC-based control algorithm to ensure adequate active and reactive power sharing between the parallel connected IBRs and the accurate tracking of the filter output filter capacitor voltage. The droop control method provides a reference current for the FCS-MPC algorithm which in turn generates the PWM signal to switch the parallel connected IBRs. We examine the behaviors of the IBRs with respect to their droop coefficients. Also, the effect of sudden changes in the load on the droop characteristics of the inverters is examined. Consequently, the key contributions in this paper are:
@@ 2.1 Stability Analysis of Finite-Set Predictive Control System
Proof.
Let us consider a lyapunov function V(x) and a Lipschitz continuous function f(x) such that:
@@ 3 Model Development
The DC power supply to each IBR is stiff and the output voltages $\mathbf{u}_{dq}$ are local and unique to each inverter Watson et al. (2019).
@@ 4.1 Droop Control development
According to Laaksonen et al. (2005), if $\delta_{ij}$ is very small for a highly inductive transmission line. That is, $\sin\delta_{ij}\approx\delta_{ij}$ and $\cos\delta_{ij}\approx 1$, where $\delta_{ij}$ is the phase shift angle, $v_{i}$ and $v_{j}$ are the output voltages of each inverter and $X_{j}$ is the inductive reactance of the transmission line. Then, equation for the output active power (P) and reactive power (Q) developed between an IBRi connected to a neighboring IBRj can be rewritten as
@@
The droop coefficients must be carefully selected or computed because they affect stability and the way loads are being shared between a set of inverter-based resources (IBRs)Bayhan et al. (2016). Thus, for the total power P to be shared among n inverters, the load distribution between them must satisfy the following:
@@ 4.2 Finite Control Set MPC Formulation
where $\textbf{y}(k)=[\mathbf{v}_{d}^{\top},\mathbf{v}_{q}^{\top},\mathbf{i}_{d}^{\top},\mathbf{i}_{q}^{\top}]^{\top}$ and $C\in\mathbb{R}^{{2n}\times{4n}}$. The reference currents $\textbf{i}_{dqref}(k)=[\textbf{i}_{dref}(k)^{\top},\textbf{i}_{qref}(k)^{\top}]^{\top}$ for the FCS-MPC controller is generated from the filter dynamic by using PI controller in the voltage control loop as shown in (19) Subramanian and B Gooi (2018).
@@ 5 Numerical Simulation Results
Figure 5(a) shows that efficient frequency synchronization and active power sharing with fewer oscillations are achieved at a steady state. Frequency deviation is minimal and reaches a steady state quickly during connection and disconnection of load at bus 3. Active power is equally shared proportional to the stiff frequency maintained in the AC grid. This shows the effectiveness of the decentralized droop-based FCS-MPC control for active power sharing. Slight frequency deviation is normal based on (9) but this can be further improved by a higher-level controller which is not considered here. Figure 5(b) shows the reactive power sharing for the output voltages of IBR1, IBR2, and IBR3. Reactive power changes with respect to changes in load and this deviation is more pronounced at bus 3 where additional load is added. Although the inverter voltages do not deviate significantly from each but the reactive is more pronounced at higher load changes. Droop-based FCS-MPC reduces chattering and oscillations in reactive power sharing and the slight deviation is due to the grid impedance mismatch Zhong (2011). And has been addressed in literature by using virtual impedance droop control Anubi and Ameli (2022); He and Li (2011). Unlike the P-f droop, the Q-V droop is local to each inverter depending on how long they are connected to each other with respect to line impedance variations. However, the reactive powers and the voltages of the IBRs reach a steady state quickly as the load connects and disconnects at bus 3 efficiently. This shows the effectiveness of FCS-MPC in the inner control loop.

### v2

@@
Decentralized Droop-based Finite-Control-Set Model Predictive Control of Inverter-based Resources in Islanded AC Microgrid
Ayobami Olajube, Koto Omiloli, Satish Vedula, and
Olugbenga Moses Anubi
Department of Electrical and Computer Engineering, the Center for Advanced Power Systems, Florida State University E-mail: {aolajube, kao23a , svedula , oanubi}@fsu.edu
1 Abstract
@@
Keywords: Droop control, finite-control-set model predictive control, active power sharing, reactive power sharing, voltage regulation, inverter-based resources, islanded microgrid.
@@ 2 Introduction
Traditional Microgrids (MGs) are a conglomeration of numerous fossil-fuel-driven power generation elements like synchronous machines (SMs). However, in recent times, an increase in penetration of inverter-based resources (IBRs) such as solar and wind generation has caused a shift in power generation trends in the existing infrastructure. The addition of energy storage elements (ESEs) further upped the choice of power allocation in the MGs in terms of the available degree of freedom. Based on their configuration and the mode of operation MGs are classified into two categories known as grid-connected and islanded modes [1]. Control of the traditional MG employs a hierarchical control structure [2], with droop-control being the lowest level followed by the power and the energy management (EM) layers. The impending integration of IBRs into the traditional MGs is giving rise to structural and operational changes in the form of transitioning away from the SMs which are ramp-limited due to their inertial properties. However, the low inertia and the fast switching properties of the IBRs pose a major challenge in the form of assessing their stability and robustness [3]. One such method that is widely embraced to restore the inertia to the grid is to introduce virtual inertia [4, 5].
@@
Regardless of the inertial properties of the SMs and the IBRs, the control objective remains unequivocal in terms of balancing the active and reactive power in the MGs while simultaneously maintaining the grid frequency around the desired operating point. Droop control is the method widely employed in addressing the control of the power and the frequency namely voltage droop and frequency droop [6]. The IBRs are operated in parallel mode due to high reliability and flexibility in terms of operation. However, the application of droop control in such a parallel setting is not suitable while sharing a nonlinear load due to the presence of harmonic currents [7]. Moreover, the frequency and amplitude deviation in the IBRs droop settings are dependent on the load scenarios [8]. Addition of virtual control loops such as adding a virtual resistance into the droop methods and decentralizing the control (consensus-based) are proposed to address the complications arising from the presence of current harmonics [9]. Although Communication-based and consensus control algorithms offer efficient voltage regulation and effective power sharing among connected IBRs, the methods suffer from the huge cost and reliability of communication links. On the other hand, droop control does not require communication between IBRs to effect power sharing and voltage regulation [10]. Multiple loops of conventional droop techniques often result in poor power-sharing due to parameter variations, especially in the current loop that generates PWM signals for the inverters [11].
@@
Thus, in this paper, we utilize FCS-MPC in the inner loop to directly generate optimal switching signals for the IBRs. Finite-control-set model predictive control (FCS-MPC) has emerged as a promising discrete-time model-based nonlinear predictive control algorithm applicable to fast-switching power electronics dominated systems. Unlike the continuous-set MPC that requires modulators and a longer horizon to generate the pulse-width modulated (PWM) signals, FCS-MPC eliminates the need for modulators and accounts for nonlinearities in the power converter dynamics [12]. It uses the horizon of h = 1, which significantly reduces the computational burden that may result from very low sampling time in high-frequency converters. To this end, we develop a decentralized droop-based FCS-MPC-based control algorithm to ensure adequate active and reactive power sharing between the parallel connected IBRs and the accurate tracking of the filter output filter capacitor voltage. The droop control method provides a reference current for the FCS-MPC algorithm which in turn generates the PWM signal to switch the parallel connected IBRs. We examine the behaviors of the IBRs with respect to their droop coefficients. Also, the effect of sudden changes in the load on the droop characteristics of the inverters is examined. Consequently, the key contributions in this paper are:
@@ 3.1 Stability Analysis of Finite-Set Predictive Control System
Proof: Let us consider a lyapunov function V(x) and a Lipschitz continuous function f(x) such that:
@@ 4 Model Development
The DC power supply to each IBR is stiff and the output voltages $\mathbf{u}_{dq}$ are local and unique to each inverter [13].
@@ 5.1 Droop Control development
According to [14], if $\delta_{ij}$ is very small for a highly inductive transmission line. That is, $\sin\delta_{ij}\approx\delta_{ij}$ and $\cos\delta_{ij}\approx 1$, where $\delta_{ij}$ is the phase shift angle, $v_{i}$ and $v_{j}$ are the output voltages of each inverter and $X_{j}$ is the inductive reactance of the transmission line. Then, equation for the output active power (P) and reactive power (Q) developed between an IBRi connected to a neighboring IBRj can be rewritten as
@@
The droop coefficients must be carefully selected or computed because they affect stability and the way loads are being shared between a set of inverter-based resources (IBRs)[12]. Thus, for the total power P to be shared among n inverters, the load distribution between them must satisfy the following:
@@ 5.2 Finite Control Set MPC Formulation
where $\textbf{y}(k)=[\mathbf{v}_{d}^{\top},\mathbf{v}_{q}^{\top},\mathbf{i}_{d}^{\top},\mathbf{i}_{q}^{\top}]^{\top}$ and $C\in\mathbb{R}^{{2n}\times{4n}}$. The reference currents $\textbf{i}_{dqref}(k)=[\textbf{i}_{dref}(k)^{\top},\textbf{i}_{qref}(k)^{\top}]^{\top}$ for the FCS-MPC controller is generated from the filter dynamic by using PI controller in the voltage control loop as shown in (19) [15].
@@ 6 Numerical Simulation Results
Figure 5(a) shows that efficient frequency synchronization and active power sharing with fewer oscillations are achieved at a steady state. Frequency deviation is minimal and reaches a steady state quickly during connection and disconnection of load at bus 3. Active power is equally shared proportional to the stiff frequency maintained in the AC grid. This shows the effectiveness of the decentralized droop-based FCS-MPC control for active power sharing. Slight frequency deviation is normal based on (9) but this can be further improved by a higher-level controller which is not considered here. Figure 5(b) shows the reactive power sharing for the output voltages of IBR1, IBR2, and IBR3. Reactive power changes with respect to changes in load and this deviation is more pronounced at bus 3 where additional load is added. Although the inverter voltages do not deviate significantly from each but the reactive is more pronounced at higher load changes. Droop-based FCS-MPC reduces chattering and oscillations in reactive power sharing and the slight deviation is due to the grid impedance mismatch [16]. And has been addressed in literature by using virtual impedance droop control [4, 17]. Unlike the P-f droop, the Q-V droop is local to each inverter depending on how long they are connected to each other with respect to line impedance variations. However, the reactive powers and the voltages of the IBRs reach a steady state quickly as the load connects and disconnects at bus 3 efficiently. This shows the effectiveness of FCS-MPC in the inner control loop.